OpenStudy (anonymous):

After its brakes are applied, a certain sports car decelerates at a constant rate of 28ft/s2. Compute the stopping distance if the car is going 88 ft/s (60 mi/h) when the brakes are applied. stopping distance =__________feet

OpenStudy (anonymous):

@sourwing help :)

OpenStudy (anonymous):

apply v= v0+at to find t (don't forget when stopping, v=0) then plug back to equation of the distance: h= h0 +v0t +1/2at^2 where h0=0